# Evaluating MCC for Low-Frequency Cyberattack Detection in Imbalanced Intrusion Detection Data

### v1

@@ Abstract.
In many real-world network environments, several types of cyberattacks occur at very low rates compared to benign traffic, making them difficult for intrusion detection systems (IDS) to detect reliably. This imbalance causes traditional evaluation metrics, such as accuracy, to often overstate model performance in these conditions, masking failures on minority attack classes that are most important in practice. In this paper, we evaluate a set of base and meta classifiers on low-traffic attacks in the CSE-CIC-IDS2018 dataset and compare their reliability in terms of accuracy and Matthews Correlation Coefficient (MCC). The results show that accuracy consistently inflates performance, while MCC provides a more accurate assessment of a classifier’s performance across both majority and minority classes. Meta-classification methods, such as LogitBoost and AdaBoost, demonstrate more effective minority class detection when measured by MCC, revealing trends that accuracy fails to capture. These findings establish the need for imbalance-aware evaluation and make MCC a more trustworthy metric for IDS research involving low-traffic cyberattacks.
@@ 1. Introduction
In this paper, we evaluate a set of base and meta-classifiers on low-traffic attacks in the CSE-CIC-IDS2018 dataset (sharafaldin2018toward). We compare performance using accuracy and the Matthews Correlation Coefficient (MCC), a metric designed to handle imbalanced data. Our results show that accuracy consistently overestimates classifier performance. In contrast, MCC provides a clearer view of strengths and weaknesses across both majority and minority attack classes, highlighting the need for IDS research to evaluate models using metrics that account for class imbalance rather than relying solely on accuracy.
@@ 2. Related Work
Although prior research has explored advanced architectures, feature engineering, and imbalance-mitigation strategies, relatively few studies analyze how evaluation metrics themselves can shape conclusions about classifier performance. This research gap motivates the present work, which compares advanced meta-classifiers and baseline models across three low-frequency attacks—Heartbleed, Web Attacks, and Infiltration—in the CSE-CIC-IDS2018 dataset. By contrasting accuracy with MCC, the study demonstrates how metric choice can alter model rankings and provide a more reliable assessment of IDS performance on rare cyberattacks.
@@ 8. Conclusion
Table 2. Classifier Performance on the Infiltration Subset of CICIDS2017
Classifier Accuracy (%) Recall (% Attack) MCC
Decision Trees
J48 99.9152 61.1 73.3
RandomForest 99.9501 72.2 85.0
RandomTree 99.9052 63.9 71.1
REPTree 99.9451 72.2 83.4
DecisionStump 99.8203 0.0 0.0
Rule-Based Learners
JRip 99.9551 83.3 87.0
OneR 99.9750 100.0 93.7
PART 99.9351 69.44 80.2
DecisionTable 99.9950 100.0 98.6
ZeroR 99.8203 0.0 0.0
Lazy / Instance-Based
IBk 99.9251 63.9 76.6
Bayesian Models
BayesNet 99.9138 80.56 22.7
NaiveBayesMultinomialText 99.8203 0.0 0.0
Meta/Ensemble Methods
Bagging 99.9301 63.9 78.2
AdaBoostM1 99.9301 61.1 78.1
ClassificationViaRegression 99.9301 94.4 94.4
AttributeSelectedClassifier 99.8852 61.1 65.8
FilteredClassifier 99.9551 86.1 87.3
IterativeClassifierOptimizer 99.9401 66.7 81.6
LogitBoost 99.9401 66.7 81.6
RandomCommittee 99.9401 72.2 81.9
MultiScheme 99.8203 0.0 0.0
CVParameterSelection 99.8203 0.0 0.0
Stacking 99.8203 0.0 0.0
Vote 99.8203 0.0 0.0
WeightedInstancesHandlerWrapper 99.8203 0.0 0.0
InputMappedClassifier 99.8203 0.0 0.0
RandomSubSpace 99.9451 72.2 83.4
Function-Based
SGDText 99.8203 0.0 0.0
VotedPerceptron 99.8203 0.0 0.0
Table 3. Classifier Performance on the Web Attack Subset of CICIDS2017
Classifier Accuracy (%) Recall (% Attack) MCC
Decision Trees
J48 99.8422 99.2 99.1
RandomForest 99.8918 99.1 99.4
RandomTree 99.7565 98.8 98.6
REPTree 99.9053 99.6 99.5
DecisionStump 94.6979 83.4 73.0
Rule-Based Learners
JRip 99.8377 99.2 99.1
OneR 99.3251 98.1 96.1
PART 99.9008 99.5 99.4
DecisionTable 98.2642 97.9 91.0
ZeroR 99.8203 0.0 0.0
Lazy / Instance-Based
IBk 99.8370 99.3 99.1
Bayesian Models
BayesNet 97.8796 97.5 89.2
NaiveBayesMultinomialText 99.8203 0.0 0.0
Meta/Ensemble Methods
Bagging 99.9594 99.7 99.6
AdaBoostM1 97.3150 83.4 84.2
ClassificationViaRegression 99.9098 99.6 99.5
AttributeSelectedClassifier 99.7701 99.1 98.7
FilteredClassifier 99.8780 99.3 99.3
IterativeClassifierOptimizer 98.8320 89.6 93.1
LogitBoost 98.8320 89.6 93.1
RandomCommittee 99.8740 99.1 99.3
MultiScheme 99.8203 0.0 0.0
CVParameterSelection 99.8203 0.0 0.0
Stacking 99.8203 0.0 0.0
Vote 99.8203 0.0 0.0
WeightedInstancesHandlerWrapper 99.8203 0.0 0.0
InputMappedClassifier 99.8203 0.0 0.0
RandomSubSpace 99.9010 99.1 99.4
Function-Based
SGDText 99.8203 0.0 0.0
VotedPerceptron 96.5239 83.7 80.6
Table 4. Classifier Performance on the Heartbleed Subset of CICIDS2017
Classifier Accuracy (%) Recall (% Attack) MCC
Decision Trees
J48 99.994 90.9 87.0
RandomForest 99.998 90.9 95.3
RandomTree 100.0 100.0 100.0
REPTree 99.996 100.0 92.0
DecisionStump 100.0 100.0 100.0
Rule-Based Learners
JRip 99.998 90.9 95.3
OneR 99.992 90.9 83.6
PART 99.996 90.9 87.0
DecisionTable 99.992 90.9 83.6
ZeroR 99.978 0.0 0.0
Lazy / Instance-Based
IBk 99.998 90.9 95.3
Bayesian Models
BayesNet 99.976 100.0 69.1
NaiveBayesMultinomialText 99.978 0.0 0.0
Meta/Ensemble Methods
Bagging 100.0 100.0 100.0
AdaBoostM1 100.0 100.0 100.0
ClassificationViaRegression 99.996 90.9 90.9
AttributeSelectedClassifier 99.996 90.9 90.9
FilteredClassifier 100.0 100.0 100.0
IterativeClassifierOptimizer 100.0 100.0 100.0
LogitBoost 100.0 100.0 100.0
RandomCommittee 99.998 90.9 95.3
MultiScheme 99.978 0.0 0.0
CVParameterSelection 99.978 0.0 0.0
Stacking 99.978 0.0 0.0
Vote 99.978 0.0 0.0
WeightedInstancesHandlerWrapper 99.978 0.0 0.0
InputMappedClassifier 99.978 0.0 0.0
RandomSubSpace 99.998 100.0 95.7
Function-Based
SGDText 99.978 0.0 0.0
VotedPerceptron 99.978 0.0 0.0

### v2

@@ Abstract.
In many real-world network environments, several types of cyberattacks occur at very low rates compared to benign traffic, making them difficult for intrusion detection systems (IDS) to detect reliably. This imbalance causes traditional evaluation metrics, such as accuracy, to often overstate model performance in these conditions, masking failures on minority attack classes that are most important in practice. In this paper, we evaluate a set of base and meta classifiers on low-traffic attacks in the CSE-CIC-IDS2017 dataset and compare their reliability in terms of accuracy and Matthews Correlation Coefficient (MCC). The results show that accuracy consistently inflates performance, while MCC provides a more accurate assessment of a classifier’s performance across both majority and minority classes. Meta-classification methods, such as LogitBoost and AdaBoost, demonstrate more effective minority class detection when measured by MCC, revealing trends that accuracy fails to capture. These findings establish the need for imbalance-aware evaluation and make MCC a more trustworthy metric for IDS research involving low-traffic cyberattacks.
@@ 1. Introduction
In this paper, we evaluate a set of base and meta-classifiers on low-traffic attacks in the CSE-CIC-IDS2017 dataset (sharafaldin2018toward). We compare performance using accuracy and the Matthews Correlation Coefficient (MCC), a metric designed to handle imbalanced data. Our results show that accuracy consistently overestimates classifier performance. In contrast, MCC provides a clearer view of strengths and weaknesses across both majority and minority attack classes, highlighting the need for IDS research to evaluate models using metrics that account for class imbalance rather than relying solely on accuracy.
@@ 2. Related Work
Although prior research has explored advanced architectures, feature engineering, and imbalance-mitigation strategies, relatively few studies analyze how evaluation metrics themselves can shape conclusions about classifier performance. This research gap motivates the present work, which compares advanced meta-classifiers and baseline models across three low-frequency attacks—Heartbleed, Web Attacks, and Infiltration—in the CSE-CIC-IDS2017 dataset. By contrasting accuracy with MCC, the study demonstrates how metric choice can alter model rankings and provide a more reliable assessment of IDS performance on rare cyberattacks.